# A Reproducible and Fair Evaluation of Partition-aware Collaborative Filtering

### v1

@@ Abstract
Similarity-based collaborative filtering (CF) models have long demonstrated strong offline performance and conceptual simplicity. However, their scalability is limited by the quadratic cost of maintaining dense item–item similarity matrices. Partitioning-based paradigms have recently emerged as an effective strategy to balance effectiveness and efficiency, allowing models to learn local similarities within coherent subgraphs while maintaining limited global context. In this work, we focus on the Fine-tuning Partition-aware Similarity Refinement (FPSR) framework, a prominent representative of this family, and its extension FPSR+. Reproducible evaluation of partition‑aware collaborative filtering remains challenging, as prior FPSR/FPSR+ reports often rely on splits of unclear provenance and omit some similarity‑based baselines, complicating fair comparison. We present a transparent, fully reproducible benchmark of FPSR and FPSR+. Based on our results, the family of FPSR models does not consistently perform at the highest level. Overall, it remains competitive, validates its design choices, and shows significant advantages in long-tail scenarios. This highlights the accuracy–coverage trade-offs resulting from partitioning, global components, and hub design. Our investigation clarifies when partition‑aware similarity modeling is most beneficial and offers actionable guidance for scalable recommender system design under reproducible protocols. Source code at https://split.to/rep_ecir.
@@ 1 Introduction
Motivation. Over the last decades, recommender systems have evolved through multiple waves of methods, from early neighborhood and latent-factor models to contemporary neural and graph-based approaches [DBLP:journals/tkde/AdomaviciusT05, MF]. Within this broader landscape, two influential strands have emerged: graph-based models, which propagate signals over user–item graphs to capture high-order relations [DBLP:conf/iclr/KipfW17, NGCF, LightGCN], and item–item similarity models, which leverage local neighborhoods to deliver simple and effective recommendations [sarwar2001, DBLP:journals/internet/LindenSY03, deshpande2004]. Each strand brings distinct strengths and liabilities: graph-based methods model multi-hop structure but typically rely on message passing and iterative training [LightGCN], whereas item–item approaches enable lightweight inference and transparent pipelines, yet face quadratic similarity computation as catalogs grow (with sparsity-inducing learners like SLIM mitigating scalability in practice) [SLIM].
@@
Against this background, partition-aware strategies have been proposed to balance accuracy and efficiency by explicitly decomposing the item graph into coherent subgraphs and learning local similarities within each part while preserving limited global context. The Fine-tuning Partition-aware Item Similarities (FPSR) framework exemplifies this direction by coupling partition-level training with a small global component, and FPSR+ augments it with hub sets that bridge partitions to enhance robustness under small or imbalanced clusters [FPSRplus]. A complementary block-aware line, instantiated by the Block-aware Item Similarity Model (BISM), regularizes a global similarity matrix so that item communities emerge as approximately block-diagonal structures during training [BISM]. Both perspectives are grounded in the empirical regularity that items in large catalogs tend to organize into cohesive communities (e.g., through co-purchase and substitution/complement networks) [DBLP:conf/sigir/McAuleyTSH15, DBLP:journals/tweb/LeskovecAH07]. The main difference is how this prior is designed: through explicit partitioning and hubs in FPSR/FPSR+ versus implicit block structure via regularization in BISM, resulting in distinct computational profiles. Given that these models leverage naturally occurring clustering yet differ fundamentally in their design choices, it is essential to investigate their performance using transparent splits, rigorous hyperparameter optimization, and metrics that capture both accuracy and beyond-accuracy dimensions.
@@
Open Issues. Despite promising reports, several issues motivate a deeper, transparent reassessment of partition-aware models. Replicability remains a concern [DBLP:journals/umuai/BelloginS21, dacrema2019, dacrema2025]: prior evaluations have relied on dataset splits that are not always public and, in some cases, validation procedures that may blur selection boundaries, complicating independent verification. Fair comparison is also open [DBLP:conf/recsys/SunY00Q0G20]: the claimed advantages of FPSR/FPSR+ over strong similarity-based baselines such as BISM warrant re-evaluation on a level playing field with consistent data splits and comparable hyperparameter search across competitors. Both FPSR/FPSR+ and BISM utilize natural clustering in user-item data, but they differ in key design choices. These differences raise questions about how each method shares information across clusters and ultimately impacts the quality of recommendations. Moreover, the capability of this family of systems to recommend long-tail items (i.e., those that receive fewer interactions), a property crucial for real-world applications [jannach2013, ge2010], has not yet been thoroughly investigated under matched protocols. FPSR/FPSR+ and BISM both leverage naturally occurring clustering, yet they differ in terms of global versus local optimization and the roles of global connectors (hub sets) and global regularization. Therefore, it is essential to systematically investigate their performance using transparent splits and metrics, with a particular focus on their effects on both the head and long tail of the catalog, a well-documented challenge in recommendation practice [park2008].
@@ 2 Related work and Background
The earliest collaborative filtering (CF) systems were memory-based neighborhood models (e.g., item-NN [deshpande2004]), which rank candidate items using pre-defined similarity functions over historical co-consumption patterns. These methods are simple and explainable, but they rely on static similarity measures and face a quadratic storage and computation bottleneck as the item catalog grows. In parallel, model-based approaches emerged, either factorizing the user–item signal into latent embeddings (e.g., MF [MF]) or learning the item–item similarity weights directly as optimization variables (e.g., SLIM [SLIM] and EASE [EASEr]). In addition to these similarity-based models, graph convolutional networks (GCNs) have gained prominence [LightGCN, UltraGCN, SVD-GCN, NGCF]. GCNs propagate user–item signals through the interaction graph to capture high-order connectivity patterns. Despite their expressive power, GCNs often have high computational overhead and slow iterative training, which limits their scalability in large-scale recommendation scenarios. This shows that data-driven similarity estimation typically surpasses fixed heuristics, while maintaining NN-like efficiency during inference.
A consistent empirical observation across learned similarity models [LorSLIM] is that items tend to form cohesive clusters, i.e., the learned similarity matrix exhibits an approximately block-diagonal structure with strong intra-block and weak inter-block relations. Interestingly, a similar phenomenon can be observed in graph-based models such as GCNs, where sampling or pruning often yields disconnected or weakly connected subgraphs that reflect natural item communities. This suggests that clustering is a unifying structural principle underlying both item-similarity and graph-based paradigms. This emergent clustering behavior had previously been treated as a by-product of learning dynamics rather than an explicit design target. BISM was the first to formalize this phenomenon, introducing it as an architectural prior rather than a side effect.
@@
FPSR further strengthens robustness in the presence of small or imbalanced partitions by introducing a hub set: a carefully selected subset of cross-partition bridge items (e.g., popular ones) whose signals are shared to stabilize local learning and recover inter-partition effects. Both FPSR and FPSR demonstrate competitive accuracy compared to strong baselines while significantly lowering computational costs and substantially reducing the model footprint through partition-wise splitting.
@@ 3 Experimental Setup
Our experimental procedure has two phases. First, we conduct a replicability study to validate the original findings. We initially ran a series of replicability tests on the datasets they used: Amazon-CDs [DBLP:conf/sigir/McAuleyTSH15], Gowalla [DBLP:conf/kdd/ChoML11], Yelp2018 [DBLP:conf/isspit/KronmuellerCHD18], and Douban [DBLP:conf/cikm/ShiZLYYW15]. We follow as closely as possible the authors’ FPSR setup: we used the original data splits and the hyperparameters reported in their public repository111https://github.com/Joinn99/FPSR. The situation is more heterogeneous for the two FPSR+ variants. For FPSR+D, the authors released an implementation limited to experiments on only one dataset, which hinders broader reproducibility. However, we have experimentally observed its hyperparameters coincide with those of FPSR, apart from the introduction of an additional fixed-value parameter. In contrast, no implementation was provided for FPSR+F variant. Based on the methodological details in the original article, we developed our own implementation, enabling us to reproduce the experimental results with reasonable effort. We contacted the authors, but unfortunately, they did not respond to us, and we are unable to address this lack of details. More details are explained in Section˜4.1.
@@
For the next phase, we adopted a new and consistent data splitting and evaluation protocol for the same four datasets that were previously mentioned. We employed a user-based hold-out splitting strategy for all datasets. For each user, 15% of their interactions were randomly sampled to create the test set. Subsequently, from the remaining interactions, another 15%for each user were sampled to form the validation set. The remaining interactions constituted the training set. For these main tests, accuracy was assessed using Recall and nDCG at . All experiments were conducted using Elliot [elliot].
@@
We compared FPSR against a set of well-established baselines: Random, MostPop, Item-kNN [deshpande2004], RP [RP3beta], EASER [EASEr], GF-CF [GFCF], LightGCN [LightGCN] and BISM. It is worth noting that in [FPSR] it lacked transparency regarding baselines hyperparameter tuning. Most baselines are established state-of-the-art models, except BISM, whose results we verified using the authors’ public code since its data split was available only for BookCrossing. Although we omit the reproducibility analysis to focus on FPSR and FPSR+, we include BISM due to its strong similarity and limited use in literature.
@@
Following the rigorous evaluation principles advocated by [dacrema2019], we took particular care to ensure a fair comparison: we performed a hyperparameter optimization for all baselines by running 20 trials based on the Tree-structured Parzen Estimator [DBLP:conf/icml/BergstraYC13] strategy, with Recall@20 as the validation metric, ensuring that each algorithm could perform at its maximum potential. This rigorous methodology allows us to robustly validate the contribution of FPSR and its variants, ensuring that any reported improvement constitutes a genuine advancement over a robust set of competitors.
@@
The complete parameters search spaces for all other baselines and all the best founded configurations are documented in our public code repository222https://split.to/rep_ecir. Due to space constraints, we report here few search spaces for the primary subject of our study, FPSR and its variants: the weight , that is the weight of global information with respect to the local information, from {0.1, 0.2, …, 0.5}; the size ratio parameter , chosen from {0.1, 0.2, …, 0.5}.
@@ 4.1 RQ1: Replicability Study
We compared in Table 2 the values published in the original paper with those calculated in our tests, accompanied by their difference. The performances are practically identical on Gowalla and Yelp2018, exhibiting a marginal performance shift on the order of while on Amazon-CDs and especially on Douban more marked discrepancies emerge. For Gowalla and Yelp2018, the correspondence is explained by the fact that the authors published the exact data splittings they used. However, it should be noted that, as can be seen from their configuration file, they also used the test set as the validation set for model selection. Conversely, for Amazon-CDs and Douban, the absence of published data splits forced us to generate them by executing the authors’ original code. This introduces stochasticity in the data splitting, a known obstacle to replicability of the experiments [DataRec]. The considerable deviations we found from the published results are a direct consequence of this.
@@ 4.2 RQ2: Fair Benchmarking
We attribute this significant reversal to our commitment to a thorough and equitable hyperparameter optimization process for all competing models. While the original authors understandably focused on tuning their novel architecture, our independent evaluation created a level playing field. BISM, a model with a distinct block-aware regularization scheme, appears to be particularly responsive to meticulous tuning. Our extensive search likely unlocked a higher performance ceiling for BISM than what was achieved in the original comparisons. This outcome serves as a critical lesson for recommender system evaluation. It suggests that the reported superiority of a new model can sometimes be an artifact of a focused tuning effort rather than an inherent architectural advantage. A strong, established baseline like BISM, when afforded the same level of rigorous optimization, can be an even more powerful competitor [dacrema2019].
@@ 5 Conclusion
In our replication study, we found that the original FPSR results are only partially reproducible. We could match the reported performance using the official data splits, but missing implementation details prevented complete verification of the FPSR+ method. This motivated us to create a benchmark for re-evaluation.
In this evaluation, we found that the block-aware BISM model can outperform the FPSR model when both are fairly optimized, contradicting the original findings. BISM achieved higher accuracy than FPSR (including FPSR+) on datasets like Amazon-CDs and Gowalla. However, FPSR remains competitive, achieving state-of-the-art results on other benchmarks, with only a small performance gap between the base FPSR and its enhanced variants. The performance gap between the base FPSR model and its enhanced variants is notable. Adding global hub items in FPSR+ improves robustness and reduces long-tail item accuracy drops, especially with small partitions or imbalanced data. However, the effectiveness of hub-selection strategies varies by dataset: FPSR+D typically excels with popular items, while FPSR+F is expected to perform better with long-tail items. Nevertheless, this is not always the case, as FPSR+F sometimes surpasses FPSR+D in head-item accuracy and vice versa for tail items.

### v3

@@ Abstract
Similarity-based collaborative filtering (CF) models have long demonstrated strong offline performance and conceptual simplicity. However, their scalability is limited by the quadratic cost of maintaining dense item–item similarity matrices. Partitioning-based paradigms have recently emerged as an effective strategy to balance effectiveness and efficiency, allowing models to learn local similarities within coherent subgraphs while maintaining limited global context. In this work, we focus on the Fine-tuning Partition-aware Similarity Refinement (FPSR) framework, a prominent representative of this family, and its extension FPSR+. Reproducible evaluation of partition‑aware collaborative filtering remains challenging, as prior FPSR/FPSR+ reports often rely on splits of unclear provenance and omit some similarity‑based baselines, complicating fair comparison. We present a transparent, fully reproducible benchmark of FPSR and FPSR+. Based on our results, the family of FPSR models does not consistently perform at the highest level. Overall, it remains competitive, validates its design choices, and shows significant advantages in long-tail scenarios. This highlights the accuracy–coverage trade-offs resulting from partitioning, global components, and hub design. Our investigation clarifies when partition‑aware similarity modeling is most beneficial and offers actionable guidance for scalable recommender system design under reproducible protocols. Source code at https://github.com/sisinflab/A-Reproducible-and-Fair-Evaluation-of-Partition-aware-Collaborative-Filtering.
@@ 1 Introduction
Motivation. Over the last decades, recommender systems have evolved through multiple waves of methods, from early neighborhood and latent-factor models to contemporary neural and graph-based approaches [1, 15]. Within this broader landscape, two influential strands have emerged: graph-based models, which propagate signals over user–item graphs to capture high-order relations [14, 31, 12], and item–item similarity models, which leverage local neighborhoods to deliver simple and effective recommendations [26, 18, 8]. Each strand brings distinct strengths and liabilities: graph-based methods model multi-hop structure but typically rely on message passing and iterative training [12], whereas item–item approaches enable lightweight inference and transparent pipelines, yet face quadratic similarity computation as catalogs grow (with sparsity-inducing learners like SLIM mitigating scalability in practice) [22].
@@
Against this background, partition-aware strategies have been proposed to balance accuracy and efficiency by explicitly decomposing the item graph into coherent subgraphs and learning local similarities within each part while preserving limited global context. The Fine-tuning Partition-aware Item Similarities (FPSR) framework exemplifies this direction by coupling partition-level training with a small global component, and FPSR+ augments it with hub sets that bridge partitions to enhance robustness under small or imbalanced clusters [32]. A complementary block-aware line, instantiated by the Block-aware Item Similarity Model (BISM), regularizes a global similarity matrix so that item communities emerge as approximately block-diagonal structures during training [5]. Both perspectives are grounded in the empirical regularity that items in large catalogs tend to organize into cohesive communities (e.g., through co-purchase and substitution/complement networks) [21, 17]. The main difference is how this prior is designed: through explicit partitioning and hubs in FPSR/FPSR+ versus implicit block structure via regularization in BISM, resulting in distinct computational profiles. Given that these models leverage naturally occurring clustering yet differ fundamentally in their design choices, it is essential to investigate their performance using transparent splits, rigorous hyperparameter optimization, and metrics that capture both accuracy and beyond-accuracy dimensions.
@@
Open Issues. Despite promising reports, several issues motivate a deeper, transparent reassessment of partition-aware models. Replicability remains a concern [3, 10, 9]: prior evaluations have relied on dataset splits that are not always public and, in some cases, validation procedures that may blur selection boundaries, complicating independent verification. Fair comparison is also open [30]: the claimed advantages of FPSR/FPSR+ over strong similarity-based baselines such as BISM warrant re-evaluation on a level playing field with consistent data splits and comparable hyperparameter search across competitors. Both FPSR/FPSR+ and BISM utilize natural clustering in user-item data, but they differ in key design choices. These differences raise questions about how each method shares information across clusters and ultimately impacts the quality of recommendations. Moreover, the capability of this family of systems to recommend long-tail items (i.e., those that receive fewer interactions), a property crucial for real-world applications [13, 11], has not yet been thoroughly investigated under matched protocols. FPSR/FPSR+ and BISM both leverage naturally occurring clustering, yet they differ in terms of global versus local optimization and the roles of global connectors (hub sets) and global regularization. Therefore, it is essential to systematically investigate their performance using transparent splits and metrics, with a particular focus on their effects on both the head and long tail of the catalog, a well-documented challenge in recommendation practice [23].
@@ 2 Related work and Background
The earliest collaborative filtering (CF) systems were memory-based neighborhood models (e.g., item-NN [8]), which rank candidate items using pre-defined similarity functions over historical co-consumption patterns. These methods are simple and explainable, but they rely on static similarity measures and face a quadratic storage and computation bottleneck as the item catalog grows. By explainable, we mean that recommendations can be justified through a small set of influential neighbor items and their similarity weights, enabling straightforward inspection of the contribution of each neighbor to the final score.
In parallel, model-based approaches emerged, either factorizing the user–item signal into latent embeddings (e.g., MF [15]) or learning the item–item similarity weights directly as optimization variables (e.g., SLIM [22] and EASE [29]). In addition to these similarity-based models, graph convolutional networks (GCNs) have gained prominence [12, 20, 25, 31]. GCNs propagate user–item signals through the interaction graph to capture high-order connectivity patterns. Despite their expressive power, GCNs often have high computational overhead and slow iterative training, which limits their scalability in large-scale recommendation scenarios. This shows that data-driven similarity estimation typically surpasses fixed heuristics, while maintaining NN-like efficiency during inference.
A consistent empirical observation across learned similarity models [6] is that items tend to form cohesive clusters, i.e., the learned similarity matrix exhibits an approximately block-diagonal structure with strong intra-block and weak inter-block relations. Interestingly, a similar phenomenon can be observed in graph-based models such as GCNs, where sampling or pruning often yields disconnected or weakly connected subgraphs that reflect natural item communities. This suggests that clustering is a unifying structural principle underlying both item-similarity and graph-based paradigms. This emergent clustering behavior had previously been treated as a by-product of learning dynamics rather than an explicit design target. BISM was the first to formalize this phenomenon, introducing it as an architectural prior rather than a side effect.
@@
FPSR further strengthens robustness in the presence of small or imbalanced partitions by introducing a hub set: a carefully selected subset of cross-partition bridge items (e.g., popular ones) whose signals are shared to stabilize local learning and recover inter-partition effects. Intuitively, when partitions become small or imbalanced, local similarity estimation becomes noisy because co-occurrence evidence is fragmented within each subgraph. The hub set acts as a shared set of anchor items across partitions, reintroducing cross-partition signal and reducing variance in the learned similarities, which improves stability under aggressive partitioning. Both FPSR and FPSR demonstrate competitive accuracy compared to strong baselines while significantly lowering computational costs and substantially reducing the model footprint through partition-wise splitting.
@@ 3 Experimental Setup
Our experimental procedure has two phases. First, we conduct a replicability study to validate the original findings. We initially ran a series of replicability tests on the datasets used in the original FPSR study: Amazon-CDs [21], Gowalla [7], Yelp2018 [16], and Douban [28]. Dataset statistics are reported in Table˜1. We follow as closely as possible the authors’ FPSR setup: we used the original data splits and the hyperparameters reported in their public repository111https://github.com/Joinn99/FPSR. The situation is more heterogeneous for the two FPSR+ variants. For FPSR+D, the authors released an implementation limited to experiments on only one dataset, which hinders broader reproducibility. However, we have experimentally observed its hyperparameters coincide with those of FPSR, apart from the introduction of an additional fixed-value parameter. In contrast, no implementation was provided for FPSR+F variant. Based on the methodological details in the original article, we developed our own implementation, enabling us to reproduce the experimental results with a standard implementation effort. We contacted the original authors; however, we did not receive a response, and we are unable to address this lack of details. More details are explained in Section˜4.1.
@@
For the next phase, we adopted a new and consistent data splitting and evaluation protocol for the same four datasets that were previously mentioned. We employed a user-based hold-out splitting strategy for all datasets. For each user, 15% of their interactions were randomly sampled to create the test set. Subsequently, from the remaining interactions, another 15% for each user were sampled to form the validation set. The remaining interactions constituted the training set. For these main tests, accuracy was assessed using Recall and nDCG at . All experiments were conducted using Elliot [2].
@@
We compared FPSR against a set of well-established baselines: Random, MostPop, Item-kNN [8], RP [24], EASER [29], GF-CF [27], LightGCN [12] and BISM. It is worth noting that Wei et al. [33] do not report details regarding baselines hyperparameter tuning. Most baselines are established state-of-the-art models, except BISM, whose results we verified using the authors’ public code since its data split was available only for BookCrossing. Although we omit the reproducibility analysis to focus on FPSR and FPSR+, we include BISM due to its strong similarity and limited use in literature.
@@
Following the rigorous evaluation principles advocated by [10], we took particular care to ensure a fair comparison: we performed a hyperparameter optimization for all baselines by running 20 trials based on the Tree-structured Parzen Estimator (TPE) [4] strategy, with Recall@20 as the validation metric, ensuring that each algorithm could perform at its maximum potential. TPE is a Bayesian optimization method that models promising hyperparameter regions by comparing configurations that yield good vs. poor validation performance, and then samples new trials from the regions expected to improve the objective. This rigorous methodology allows us to robustly validate the contribution of FPSR and its variants, ensuring that any reported improvement constitutes a genuine advancement over a robust set of competitors.
@@
The complete parameters search spaces for all other baselines and all the best founded configurations are documented in our public code repository222https://github.com/sisinflab/A-Reproducible-and-Fair-Evaluation-of-Partition-aware-Collaborative-Filtering. Due to space constraints, we report here few search spaces for the primary subject of our study, FPSR and its variants: the weight , that is the weight of global information with respect to the local information, from {0.1, 0.2, , 0.5}; the size ratio parameter , chosen from {0.1, 0.2, , 0.5}.
@@ 4.1 RQ1: Replicability Study
We compared in Table 2 the values published in the original paper with those calculated in our tests, accompanied by their difference. The performances are practically identical on Gowalla and Yelp2018, exhibiting a marginal performance shift on the order of while on Amazon-CDs and especially on Douban more marked discrepancies emerge. For Gowalla and Yelp2018, the correspondence is explained by the fact that the authors published the exact data splittings they used. However, it should be noted that, based on the configuration file provided in the original FPSR repository, it appears that the test split was also used for model selection. We note that this is an inference from the released configuration and may not fully reflect all experimental choices described in the paper. Conversely, for Amazon-CDs and Douban, the absence of published data splits forced us to generate them by executing the authors’ original code. This introduces stochasticity in the data splitting, a known obstacle to replicability of the experiments [19]. The considerable deviations we observed from the published results are a direct consequence of this.
@@ 4.2 RQ2: Fair Benchmarking
We attribute this significant reversal to our commitment to a thorough and equitable hyperparameter optimization process for all competing models. While the original authors understandably focused on tuning their novel architecture, our independent evaluation created a level playing field. BISM, a model with a distinct block-aware regularization scheme, appears to be particularly responsive to meticulous tuning. Our extensive search likely may have revealed a higher performance ceiling for BISM than what was achieved in the original comparisons. This outcome serves as a critical lesson for recommender system evaluation. It suggests that the reported superiority of a new model can sometimes be an artifact of a focused tuning effort rather than an inherent architectural advantage. A strong, established baseline like BISM, when afforded the same level of rigorous optimization, can be an even more powerful competitor [10].
@@ 5 Conclusion
In our replication study, we observed that the original FPSR results are only partially reproducible. We could match the reported performance using the official data splits, but missing implementation details prevented complete verification of the FPSR+ method. This motivated us to create a benchmark for re-evaluation.
In this evaluation, we observed that the block-aware BISM model can outperform the FPSR model when both are fairly optimized, contradicting the original findings. BISM achieved higher accuracy than FPSR (including FPSR+) on datasets like Amazon-CDs and Gowalla. However, FPSR remains competitive, achieving state-of-the-art results on other benchmarks, with only a small performance gap between the base FPSR and its enhanced variants. The performance gap between the base FPSR model and its enhanced variants is notable. Adding global hub items in FPSR+ improves robustness and reduces long-tail item accuracy drops, especially with small partitions or imbalanced data. However, the effectiveness of hub-selection strategies varies by dataset: FPSR typically excels with popular items, while FPSR is expected to perform better with long-tail items. Nevertheless, this is not always the case, as FPSR sometimes surpasses FPSR in head-item accuracy and vice versa for tail items.
@@
Ack
This work was partially supported by the following projects: P+ARTS - Partnership for Artistic Research in Technology and Sustainability, funded by the European Union Next- GenerationEU (NRRP – M4C1, Investment 3.4, INTAFAM00037, CUP: G43C24000640006).